OpenStudy (anonymous):

Solve: 3 over 4 = 21 over 2 + d

OpenStudy (anonymous):

3/4=21/(2+d)

OpenStudy (anonymous):

So what you would want to do is get the d+2 into the numerator.

OpenStudy (anonymous):

So that it equals 3(2+d)/4=21

OpenStudy (anonymous):

Multiply by the reciprocal (4/3) to get the 2+d by itself so that it is: 2+d=(21*3)/4

OpenStudy (anonymous):

subtract 2

OpenStudy (anonymous):

then d should equal (21*3)/4-2

OpenStudy (anonymous):

THank you for explaing this to me im cyber school and im doing summer cyber so this is alot of help thank you

OpenStudy (anonymous):

order of operations (63/4)-2

OpenStudy (anonymous):

No prob

OpenStudy (anonymous):

okay i understand this a little better now thanks! :)

OpenStudy (anonymous):

this is the right answer

OpenStudy (anonymous):

should be 26

OpenStudy (anonymous):

yeah i saw where you went wrong and i got 26 than haha its okay! :)

OpenStudy (anonymous):

lol even old dogs need to refresh on old tricks, anyways sorry about that. If you don't quite understand it then the file is linked